# Emergent Abilities of Large Language Models under Continued Pretraining for Language Adaptation

### v2

@@ 2 Experimental Setup
For Basque, we use the Latxa corpus (Etxaniz et al., 2024), which consists of 4.7B tokens of high-quality Basque text. For Arabic and Indonesian, we randomly sample documents from their respective CulturaX corpus (Nguyen et al., 2023). We ensure all languages have token count parity (4.5$\sim$4.7B tokens per language). When including English in the CPT mixture, we use a random sample of 500k English documents from the Pile (Gao et al., 2020). For all languages, the English data accounts for 20% of the total CPT tokens.
@@ 3 Copain: A Language-Agnostic ICL Benchmark
[…]
@@ 5.2 EMA of Model Parameters
$\theta_{t}=\begin{cases}\theta^{\prime}_{t}&\text{if }t\leq 0\lor t\bmod\eta\neq 0\\ \alpha\theta_{t-\eta}+\left(1-\alpha\right)\theta^{\prime}_{t}&\text{otherwise}\\ \end{cases}$
where $\theta^{\prime}_{t}$ and $\theta_{t}$ denote the model parameters at step $t$ before and after EMA is applied, respectively, and $\alpha$ denotes the decay rate, which we set to 0.92 in all of our experiments. Unless otherwise indicated, we use $\eta=1$ for Basque and Indonesian, and $\eta=10$ for Arabic.
@@ Acknowledgments
This work has been partially supported by the Basque Government (Research group funding IT1805-22). Ahmed Elhady holds a PhD grant supported by the Basque Government (IKER-GAITU project). We also thank Aimar Zabala from the HiTZ Centre for his valuable inputs in some experiments.
@@ Appendix D Multiple Choice Prompting Computation
STEM Humanities Language Social Science Local Culture

### v3

@@ 2 Experimental Setup
For Basque, we use the Latxa corpus (Etxaniz et al., 2024), which consists of 4.7B tokens of high-quality Basque text. For Arabic and Indonesian, we randomly sample documents from their respective CulturaX corpus (Nguyen et al., 2023). We ensure all languages have token count parity (4.54.7B tokens per language). When including English in the CPT mixture, we use a random sample of 500k English documents from the Pile (Gao et al., 2020). For all languages, the English data accounts for 20% of the total CPT tokens.
@@ 3 Copain: A Language-Agnostic ICL Benchmark
Max. integer in list 85, 24, 63 85
29, 47, 79 79
59, 77, 41 77
19, 81, 88
Min. integer in list 85, 24, 63 24
29, 47, 79 29
59, 77, 41 41
19, 81, 88
Median integer in list 85, 24, 63 63
29, 47, 79 47
59, 77, 41 59
19, 81, 88
Even integer among odd list 21, 71, 68, 95 68
25, 35, 58, 83 58
92, 71, 61, 29 92
97, 66, 1, 3
Odd integer among even list 24, 76, 60, 51 51
83, 52, 22, 52 83
32, 68, 10, 79 79
64, 87, 0, 28
First character alphabetically w, y, a a
b, m, k b
v, h, p h
y, e, p
Last character alphabetically w, y, a y
b, m, k m
v, h, p v
y, e, p
@@ 5.2 EMA of Model Parameters
[…]
where $\prime\theta^{\prime}_{t}$ and $\theta_{t}$ denote the model parameters at step $t$ before and after EMA is applied, respectively, and $\alpha$ denotes the decay rate, which we set to 0.92 in all of our experiments. Unless otherwise indicated, we use $\eta=1$ for Basque and Indonesian, and $\eta=10$ for Arabic.
@@ Acknowledgments
This work has been partially supported by the Basque Government (Research group funding IT1805-22). Ahmed Elhady holds a PhD grant supported by the Basque Government (IKER-GAITU project). We also thank Aimar Zabala from the HiTZ Centre for his valuable inputs in some experiments. The models were trained on the Leonardo supercomputer at CINECA under the EuroHPC Joint Undertaking, project EHPC-EXT-2024E01-042.
@@ Appendix D Multiple Choice Prompting Computation
STEM Humanities Language
Social
Science
Local
Culture